OpenStudy (anonymous):

Sally has only nickels and dimes in her money box. She knows that she has less than $20 in the box. Let x represent the number of nickels in the box and y represent the number of dimes in the box. Which of the following statements best describes the steps to graph the solution to the inequality in x and y?

OpenStudy (anonymous):

ANSWER CHOICES Draw a dashed line to represent the graph of 5x + 10y = 2000 and shade the portion above the line for positive values of x and y. Draw a dashed line to represent the graph of 10x + 5y = 2000 and shade the portion above the line for positive values of x and y. Draw a dashed line to represent the graph of 5x + 10y = 2000 and shade the portion below the line for positive values of x and y. Draw a dashed line to represent the graph of 10x – 5y = 2000 and shade the portion below the line for positive values of x and y.

OpenStudy (anonymous):

@ranga

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (solomonzelman):

Which choices can you eliminate right away, any ideas?

OpenStudy (anonymous):

@SolomonZelman no

OpenStudy (ranga):

x nickels is worth 5x cents. y dimes is worth 10y cents. Together they are worth 5x + 10y cents. The total is less than $20 or 2000 cents. So 5x + 10y < 2000

OpenStudy (anonymous):

but the answer choices arent inequalities

OpenStudy (ranga):

Since the inequality has "less than" and not "less than or equal to", the line should be a dashed line to represent the inequality 5x + 10y < 2000. And since it is a "less than" and not "greater than" the region below the line should be shaded.

OpenStudy (anonymous):

so its c

OpenStudy (ranga):

yes.

OpenStudy (anonymous):

@ranga thank you, can you help me with another question